# Diffusion MRI with Machine Learning

### v1

@@ II-A Micro-structure mapping
  • Methodologically, these works have been dominated by neural networks, although other models such as random forests [25, 21, 43] and Bayesian estimation methods [54, 55] have also been used. The first work to propose deep neural networks for this purpose was q-space deep learning [26]. It reported accurate estimation of diffusion kurtosis imaging measures (DKI [56]) using 12 measurements and the Neurite Orientation Dispersion and Density Imaging model (NODDI [57]) using 8 measurements, thereby leading to a twelve-fold reduction in scan time. Many studies have followed [26] to develop neural network models for micro-structure mapping from dMRI signal. This should not be surprising given that deep neural networks have emerged as the best-performing models in many similar applications [58]. Most non-neural-network studies for micro-structure estimation with machine learning either predate the recent surge of deep learning or do not include a rigorous comparison with them.
  • Most studies have adopted an end-to-end training approach, where the input dMRI measurements are mapped to the target parameter of interest using a single (albeit deep) neural network. However, there are notable exceptions that are instructive. For DKI, Kerkela et al. [15] used the neural network predictions as input to a regularized non-linear least-squares method to compute the final values. This approach improved the estimation robustness of the standard method and reduced the probability of predicting implausible values. Other studies have also proposed to combine neural network estimators with conventional methods. For example for computing the NODDI parameters under very low signal to noise ratio (SNR), Gong et al. used neural networks to obtain good initial estimates, which were then used to obtain more accurate estimates using maximum likelihood estimation (MLE) [53]. The justification for this approach is that when SNR is low, MLE is prone to erroneous predictions because of its vulnerability to local minima. The predictions of the neural networks, on the other hand, although very fast, can suffer from small but significant biases. Hence, one can use the neural network to compute a good initial estimate, which can then be refined using an unbiased MLE estimator. A similar approach was used in [59], where a shallow neural net was used to estimate the isotropic volume fraction in the NODDI model, which was the used in a MLE formulation to compute the full NODDI model from single-shell data. For diffusion tensor estimation from six measurements, Liu et al. argue that directly mapping the measurements to the tensor is not optimal [60]. Given the measurements along six gradient directions, they employ a neural network to estimate the signal along additional directions. These measurements are subsequently used by a second network to estimate the diffusion tensor. Ye et al. [61, 62, 63] have proposed a series of methods based on sparse representations and deep neural networks. In one implementation, their method is a two-stage pipeline: the first stage uses an LSTM model to compute a sparse decomposition of the signal in a dictionary. A multilayer perceptron (MLP) computes the desired micro-structure indices based on the sparse representation coefficients. This work has been in part inspired by the widely-used AMICO method [52], which decouples the estimation of NODDI and ActiveX [49] models into linear problems and solves them using sparsity-based optimization. Ye et al.’s experiments show that this method can estimate NODDI parameters with higher accuracy than standard estimation techniques. This method was further improved with the help of a separable spatial-angular dictionary for signal representation [64].
  • The most common training approach has been supervised training, where the loss function to be minimized is computed based on the difference between the predicted and ground truth parameter values. However, there have been important and instructive exceptions. Kaandorp propose to train a neural network model in an unsupervised manner by using the predicted tissue parameters to generate the corresponding dMRI signal and optimizing the network weights to minimize the difference between the predicted and measured data [34, 65]. They found that predictions of unsupervised methods had higher variability whereas predictions of supervised methods had a strong bias towards the mean of training data and were deceptively smooth [65]. Epstein et al. have analyzed the bias-variance trade-off in supervised and self-supervised micro-structure estimation methods [66]. They argue that the reason for the lower bias of self-supervised methods is because they are based on the same optimization objective as MLE, while the high bias in supervised method is because they deviate from this objective and are based on a poor choice of training target. They show that using an MLE-computed parameter value as the estimation target can reduce the estimation bias of supervised methods.
  • Overall, the results of the published studies suggest that machine learning methods may be capable of achieving higher estimation accuracy than conventional methods (Tables I, II and III). They may also be able to reduce the required number of measurements. Some works have found that estimation accuracy of deep learning methods may be less affected by sub-optimal acquisition protocols than numerical optimization methods [36]. There are, however, studies that have reported contrary results that deserve careful consideration. For example, one study has reported that neural network and random forest models trained on simulated data may produce less accurate and more biased results compared with traditional nonlinear least-squares estimation methods [67]. Another study has shown that deep learning methods can achieve good estimation accuracy when the optimization landscape is well-behaved but that they have poor performance when the optimization landscape is degenerate such as for multi-compartment models with even two compartments [68, 54].
  • It has also been reported that machine learning methods are less sensitive to measurement noise and other imperfections [69, 40]. As an example, using data from healthy subjects with voluntary head motion as well as attention deficit hyperactivity disorder (ADHD) patients with non-voluntary motion, it was shown that a deep learning model computed diffusion tensor imaging (DTI) and DKI parameters that were less sensitive to motion and comparable to computations with motion-free data [40].
  • It has been shown that models that leverage spatial correlations in a neighborhood (e.g., an image patch) around the voxel of interest can lead to more accurate estimation [70, 38, 71, 72]. Most earlier studies used CNNs to learn the spatial correlations. More recent studies haves relied on attention-based neural networks that are considered to be more effective in learning correlations [73, 74, 75]. Zheng et al. argue for the superiority of transformer architectures for micro-structure mapping in dMRI [73]. In order to enable effective implementation of these architectures with small datasets, they propose to introduce additional computational modules. Specifically, they use a sparse representation stage to compute the signal embeddings in a dictionary using unfolded iterative hard thresholding.
  • Most commonly, target micro-structure values for the training data are computed by applying a standard estimation method on dMRI measurements. This is often justified by using densely-acquired high-SNR measurements to improve the accuracy of this computation. Then, the machine learning model is trained with reduced measurements to match the estimation accuracy of the standard method. However, this approach inevitably inherits some of the limitations of the standard method that is used to compute the training target. Very little attention has been paid to this issue [45]. One study proposed to inspect the results computed by the standard fitting method and remove the voxels that contain implausible values [15].
  • Most studies have evaluated the new methods only in terms of estimation accuracy measures such as root mean square of the error (RMSE). However, several studies have investigated the downstream use of the estimated micro-structure parameters. For example, Gibbons et al. showed that post‐stroke outcome prediction was the same for micro-structure maps estimated with standard methods and those estimated with a neural network with 10-fold fewer measurements [70]. Aliotta et al. observed that deep learning-estimated fractional anisotropy (FA) maps resulted in more accurate delineation of the brain tumor boundaries in glioblastoma multiforme patients than with conventional methods [76], while Kaandorp et al. showed that intravoxel incoherent motion (IVIM) parameters computed by a deep learning method better predicted the chemoradiotherapy response of pancreatic cancer patients than a standard method [34]. Another work assessed the micro-structure maps estimated by several deep learning methods in terms of their effectiveness in studying the impact of migraine on the brain white matter [77]. Specifically, FA and mean diffusivity (MD) values computed by different methods were compared using the widely used Tract-Based Spatial Statistics (TBSS) [78]. It was observed that, compared with a standard estimation technique, deep learning methods improved the true positive rate but also increased the false positive rate.
  • In some applications, obtaining reliable in-vivo training data may be hard or impossible. Two examples are cardiac DTI [32] and fetal imaging [74]. For cardiac DTI, Weine et al. [32] developed a parameterized pipeline to synthesize training data with realistic spatial correlations and cardiac motion. For fetal brain DTI, another work proposed a pipeline that synergistically combined data from preterm neonates and fetuses to synthesize realistic fetal dMRI data [74, 79]. Both studies reported good results on independent real test data. Using synthetic data has the added advantage that it can incorporate a much wider range of parameters and factors such as noise and motion than can be available in any in-vivo dataset.
@@
  • For estimating the diffusion tensor, which is the most commonly used model in dMRI, several works have reported good reconstructions from the theoretical minimum of six measurements (Table I). DeepDTI [28] uses six diffusion-weighted measurements and a non-weighted (b0) measurement, T1-weighted, and T2-weighted images as input and computes the residuals with respect to high-quality measurements. The estimated high-quality measurements are then used to estimate the diffusion tensor with a standard estimation technique. Detailed analysis, including local assessment of DTI-derived parameters, tract-based analysis, and tractography show that DeepDTI achieves accurate diffusion tensor estimation while reducing the number of measurements by a factor of at least 3-5. DeepDTI is based on a 3D Fully Convolutional Network (FCN) and requires that the six diffusion-weighted measurements be acquired along the optimal directions proposed in [81]. SuperDTI, proposed by Li et al. [38], also used an FCN for direct estimation of diffusion tensor. It achieved accurate FA computation and tractography with six measurements and accurate MD computation with three measurements. Similar results were reported in [76, 72]. Using transformer networks, another study made a similar claim and further verified the claim using tractography evaluations and quantitative structural connectivity assessments [33]. DiffNet [82] estimated FA and MD with 3-20 diffusion-weighted measurements and showed that the estimation accuracy was still higher than standard methods. DiffNet-estimated FA was more accurate for segmenting brain tumor volume, while also reducing the scan time.
  • In addition to the DTI, DKI, and NODDI parameters, machine learning methods have also been applied to compute other parameters. Examples include myelin water fraction [69, 21], water residence time in white matter (related to axonal permeability) [25, 83, 43], axonal radius [25, 21], spherical mean technique (SMT) [61, 67, 71], ensemble average propagator (EAP) models [84, 61], and relaxation-diffusion models of white matter micro-structure [36, 85, 68]. Parik et al. developed a novel magnetic resonance fingerprinting sequence to simultaneously estimate T1, T2, and diffusion tensor parameters. Instead of the standard dictionary matching approach, they used a deep neural network to estimate the target parameters and achieved good results in healthy brains and brains of MS patients [86]. Moreover, a number of works have attempted to use deep neural networks to estimate the IVIM model [65, 35, 87, 34, 88, 89, 73, 66]. One study showed that an MLP could achieve results that were comparable with or more accurate than least-squares estimation and Bayesian techniques [35]. Another study found that a neural network could outperform the state of the art estimation methods for computing the parameters of a combined IVIM-kurtosis imaging model [87]. It also showed that the neural network method made fewer predictions that were outside the range of plausible parameter values. One study showed that a neural network, trained using an unsupervised approach, produced more consistent and more accurate estimations than standard methods and it better predicted the chemoradiotherapy response of pancreatic cancer patients [34].
@@
TABLE II: A summary of recent machine learning techniques for estimating the parameters of the NODDI model.
 Reference
Summary
 [59]
@@
A dictionary learning method is first used to estimate the isotropic volume fraction. Subsequently, a neural network is used to compute the remaining NODDI parameters. Results show that this method can estimate the NODDI model from a single-shell acquisition with an accuracy that is comparable with estimations from multi-shell data.
[61] [64]
These works employed such models as LSTMs and MLPs. They reported more than twice reduction in estimation error compared with optimization methods.
[70]
This work simultaneously estimated NODDI and generalized fractional anisotropy. The trained model worked well on data from stroke patients.
[92]
New loss functions are introduced to encourage accurate prediction of the dMRI signal as well as low micro-structure estimation error. The effectiveness of the new loss functions are demonstrated, but no comparison with standard estimation methods are reported.
[73]
This work makes use of sparsity-based representation of the dMRI signal, computed using an unfolded iterative hard thresholding method, and a transformer network. It achieves an eleven-fold reduction in the required number of measurements.
[63]
@@
This method is based on representing the dMRI signal in a dictionary-matching framework similar to AMICO [52]. An MLP-type network is used to compute the NODDI parameters from the signal representations. The method achieves more than twice lower estimation error and faster computation compared with standard techniques.
[26]
@@
Using a voxel-wise estimation with an MLP, this work reported a twelve-fold reduction in the required number of measurements to achieve the same level of accuracy as standard methods. As few as 8 measurements were sufficient for accurate estimation.
[53]
@@
The focus of this work is on very low SNR scenarios such as spinal cord imaging. An MLP is used to estimate good initial values for the NODDI parameters. Refined estimates are subsequently computed using a maximum likelihood estimation technique.
[93]
@@
This method follows a super-resolution framework. Specifically, the low-resolution dMRI signal volumes are used to compute high-resolution micro-structure maps. The method itself has two stages. The first stage is a learned sparse representation of the dMRI signal with 1D convolutions. The second stage is a CNN that performs the computation. Evaluations have shown that this method outperforms conventional and machine learning methods.
[94]
The main contribution of this work is to develop a pre-training strategy that allows the model to leverage a high-quality auxiliary dataset. A sparsity-based method is used to interpolate the auxiliary dataset to match the acquisition scheme of the target dataset.
[54]
This is a Bayesian estimation method, where a population-informed prior is computed from a cohort of healthy subjects and included in the estimation for test subjects. Results show that the use of the population-based prior helps alleviate the ill-posedness of the estimation problem and significantly improves the estimation accuracy and robustness.
[82]
@@
This is a CNN-based method that achieves low reconstruction error while reducing the computation time by three orders of magnitude.
[71]
This work has trained deep learning architectures to jointly estimate several micro-structural indices including NODDI parameters. Results show that, compared with standard model fitting procedures, all architectures investigated can lead to substantial reduction in the amount of dMRI data needed for the computation and significantly faster computations.
TABLE III: A summary of recent machine learning methods for DKI estimation.
 Reference
Summary
 [26]
@@
This work has reported accurate estimation of DKI parameters from 12 measurements. The model is a voxel-wise MLP, which is tested in multiple external datasets.
[95]
@@
Hierarchical CNNs are used to estimate DTI and DKI models jointly. Results show that DKI parameters can be estimated with 8 measurements.
[96]
@@
This work has trained an MLP using synthetic data, where only four diffusion-weighted measurements are used for DKI estimation. Tests on synthetic and real data show that high estimation accuracy is achieved. This work also demonstrates the importance of noise level matching between training and test datasets.
[92]
This work uses an MLP to predict new q-space shells and to improve the angular resolution of q-space data. The enhanced data are then used to estimate the DKI parameters with a second MLP.
[15]
This work has proposed to used an MLP to estimate initial DKI values, which are then used within a regularized least-squares procedure to compute the final values.
[71]
One-, two-, and three-dimensional neural networks are used to estimate the DKI model. These models achieve better estimation accuracy than conventional model fitting method. However, the deep learning methods show some systematic biases, such as underestimation of radial kurtosis.
[97]
An MLP is trained to predict the diffusion kurtosis from synthetic dMRI data. Results show that when training a deep learning model with synthetic data, using multiple levels of noise lead to better results on clinical test data than using a single noise level.
[41]
@@
The proposed method is a hierarchical CNN combined with a motion detection and rejection technique. The deep learning model is combined with a motion detection and rejection scheme. Experiments, including on data from children with attention deficit hyperactivity disorder, show that the deep learning model is highly robust to head motion can can estimate the DKI parameters from eight measurements. It works well in the presence of severe motion, when up to 90% of the data is motion-corrupted.
II-B Fiber orientation density estimation

### v2

@@ IV-D2 Method design
Methodologically, these works have been dominated by neural networks, although other models such as random forests [45, 42, 60] and Bayesian estimation methods [153, 154] have also been used. The first work to propose deep neural networks for this purpose was q-space deep learning [16]. It reported accurate estimation of DKI parameters with 12 measurements and NODDI parameters with 8 measurements, resulting in a twelve-fold reduction in scan time. Many studies have followed [16] to develop neural network models for tissue microstructure mapping. This should not be surprising given that deep neural networks have emerged as the best-performing regression models in many applications [155]. Most non-neural-network methods either predate the recent surge of deep learning or do not include a rigorous comparison with neural networks.
@@
IV-D3 Training procedures and data
Most studies have adopted an end-to-end training approach, where the input dMRI measurements are mapped to the target parameter of interest using a single (albeit deep) neural network. However, there are notable exceptions that are instructive. For DKI, Kerkela et al. [33] used the neural network predictions as input to a regularized non-linear least-squares method to compute the final values. This approach improved the estimation robustness of the standard method and reduced the probability of predicting implausible values. Other studies have also proposed to combine neural network estimators with conventional methods. For example for computing the NODDI parameters under very low signal to noise ratio (SNR), Gong et al. used neural networks to obtain good initial estimates, which were then used to obtain more accurate estimates using maximum likelihood estimation (MLE) [147]. The justification for this approach is that when SNR is low, MLE is prone to erroneous predictions because of its vulnerability to local minima. The predictions of the neural networks, on the other hand, although very fast, can suffer from small but significant biases. Hence, one can use the neural network to compute a good initial estimate, which can then be refined using an unbiased MLE estimator. A similar approach was used in [157], where a shallow neural net was used to estimate the isotropic volume fraction in the NODDI model, which was the used in a MLE formulation to compute the complete NODDI model from single-shell data.
For diffusion tensor estimation from six measurements, Liu et al. argue that directly mapping the measurements to the tensor is not optimal [151]. Given the measurements along six gradient directions, they employ a neural network to estimate the signal along additional directions. These measurements are subsequently used by a second network to estimate the diffusion tensor. Ye et al. [158, 159] have proposed methods based on sparse representations and deep neural networks. In one implementation, their method is a two-stage pipeline: the first stage uses an LSTM model to compute a sparse decomposition of the signal in a dictionary. An MLP computes the desired microstructure indices based on the sparse representation coefficients. This work has been in part inspired by the widely-used AMICO method [146], which decouples the estimation of NODDI and ActiveX [143] models into linear problems and solves them using sparsity-based optimization. Ye et al.’s experiments show that this method can estimate NODDI parameters with higher accuracy than standard estimation techniques. This method was further improved with the help of a separable spatial-angular dictionary for signal representation [160].
@@ IV-D3 Training procedures and data
TABLE III: A summary of recent machine learning techniques for estimating the parameters of the NODDI model.
  Method Summary of methodology and results
@@
  DLpN [157] A dictionary learning method is first used to estimate the isotropic volume fraction. Subsequently, a neural network computes the remaining NODDI parameters. This method estimates the NODDI parameters from a single-shell (b=2000) acquisition as accurately as conventional methods with multi-shell data. The method is also validated on clinical data.
Deep sparse representation methods [158] [160] These related works compute sparse representation of the signal in dictionaries. Sparse representation coefficients are then used by neural networks to predict the microstructure indices. Using only HCP data, more than twice reduction in estimation error compared with optimization methods are reported.
@@
MEDN [159] This method is based on representing the dMRI signal in a dictionary-matching framework inspired by AMICO [146] and related to the above two methods [158, 160]. An MLP-type network is used to compute the NODDI parameters from the signal representations. The method achieves more than twice lower estimation error and faster computation compared with standard techniques.
Gibbons et al. [161] This work simultaneously estimated NODDI and generalized FA. The trained model worked well on data from healthy individuals and stroke patients. The method is a basic 2D FCN.
Chen at al. [95] This work used MLP models to predict unmeasured dMRI signal and microstructure indices including NODDI parameters. New loss functions are introduced to encourage accurate prediction of the dMRI signal and microstructure indices. Using only HCP-style data, the effectiveness of the new loss functions are demonstrated. No comparisons with standard estimation methods are reported.
METSC [162] This work makes use of sparsity-based representation of the dMRI signal, computed using an unfolded iterative hard thresholding method, and a transformer network. It achieves an eleven-fold reduction in the required number of measurements compared with conventional methods and reduces the estimation error compared with other learning-based methods.
@@
q-space deep learning [16] Using a voxel-wise estimation with an MLP, this work reported a twelve-fold reduction in the required number of measurements to achieve the same level of accuracy as standard methods. As few as 8 measurements were sufficient for accurate estimation. Four different datasets including one from MS patients are used.
@@
Machine learning-informed estimation [147] The focus of this work is on very low SNR scenarios such as spinal cord imaging. An MLP is used to estimate good initial values for the NODDI parameters. Refined estimates are subsequently computed using a maximum likelihood estimation technique.
@@
SR-q-DL [105] This method follows a super-resolution framework. Specifically, the low-resolution dMRI signal volumes are used to compute high-resolution tissue microstructure maps. The method itself has two stages. The first stage learns a sparse representation of the dMRI signal with 1D convolutions. The second stage is a CNN that performs the computation. Evaluations have shown that this method outperforms conventional and machine learning methods. However, only HCP data is used to develop and validate the method.
@@
DIFFnet [149] This is a CNN-based method that achieves low reconstruction error while reducing the computation time by three orders of magnitude.
Population‐based Bayesian regularization [153] This is a Bayesian estimation method, where a population-informed prior is computed from a cohort of healthy subjects and included in the estimation for test subjects. Results show that the use of the population-based prior helps alleviate the ill-posedness of the estimation problem and significantly improves the estimation accuracy and robustness. The model was developed using HCP data and validated on an independent dataset.
The most common training approach has been supervised training, where the loss function to be minimized is computed based on the difference between the predicted and ground truth parameter values. However, there have been important and instructive exceptions. Kaandorp propose to train a neural network model in an unsupervised manner by using the predicted tissue parameters to predict the corresponding dMRI signal and optimizing the network weights to minimize the difference between the predicted and measured signal [52, 163]. They found that predictions of unsupervised methods had higher variability whereas predictions of supervised methods had a strong bias towards the mean of training data and were deceptively smooth [163]. Epstein et al. have analyzed the bias-variance trade-off in supervised and self-supervised microstructure estimation methods [164]. They argue that the reason for the lower bias of self-supervised methods is because they are based on the same optimization objective as MLE, while the high bias in supervised method is because they deviate from this objective and are based on a poor choice of training target. They show that using an MLE-computed parameter value as the estimation target can reduce the estimation bias of supervised methods.
Most commonly, target microstructure values for the training data are computed by applying a standard estimation method on dMRI measurements. This is often justified by using densely-acquired high-SNR measurements to improve the accuracy of this computation. Then, the machine learning model is trained with downsampled measurements to match the estimation accuracy of the standard method. However, this approach inevitably inherits some of the limitations of the standard method that is used to compute the prediction target. Very little attention has been paid to this issue [62]. One study proposed to inspect the results computed by the standard fitting method and remove the voxels that contain implausible values [33].
In some applications, obtaining reliable in-vivo training data may be hard or impossible. Two examples are cardiac DTI [50] and fetal imaging [22]. For cardiac DTI, Weine et al. [50] developed a parameterized pipeline to synthesize training data with realistic spatial correlations and cardiac motion. For fetal brain DTI, another work proposed a pipeline that synergistically combined data from preterm neonates and fetuses to synthesize realistic fetal dMRI data [22]. Both studies reported good results on independent real test data. Using synthetic data has the added advantage that it can incorporate a much wider range of parameters and factors such as noise and motion than can be available in any in-vivo dataset.
TABLE IV: A summary of recent machine learning methods for DKI estimation.
  Method Summary of methodology and results
  Separable dictionaries and deep learning [165] A comparison of different deep learning methods showed that a technique that utilized separable dictionaries [160] achieved the best results for jointly estimating DTI, DKI, and NODDI parameters. Only HCP data were used. Moreover, the deep learning methods show some systematic biases, such as underestimation of radial kurtosis.
@@
q-space deep learning [16] This work has reported accurate estimation of DKI parameters from 12 measurements. The model is a voxel-wise MLP, which is tested in multiple external datasets.
@@
Li et al. [166] The method estimates DTI and DKI metrics jointly. Results show that DKI parameters can be estimated with 8 measurements.
@@
Masutani [167] This work trained an MLP using synthetic data. Only four diffusion-weighted measurements were used for DKI estimation. Tests on synthetic and real data showed that high estimation accuracy was achieved. This work also demonstrated the importance of noise level matching between training and test datasets.
@@
Hierarchical CNNs [58] The deep learning model is a hierarchical CNN that is combined with a motion detection and rejection technique. Experiments, including data from children with attention deficit hyperactivity disorder, show that the new method is robust to head motion and can estimate the DKI parameters from eight measurements. It works well in the presence of severe motion, when up to 90% of the data are motion-corrupted.
MLP + least squares [33] This work uses an MLP to estimate initial DKI parameter values, which are then used within a regularized least-squares procedure to compute the final values. This approach avoids implausible predictions and improves model robustness compared with existing linear and non-linear optimization methods. The data includes scans of 10 healthy volunteers.
IV-D4 Evaluations and main results
Overall, the results of the published studies suggest that machine learning methods may be capable of achieving higher estimation accuracy than conventional methods (Tables II, III and IV). They may also be able to reduce the required number of measurements. Some works have found that estimation accuracy of deep learning methods may be less affected by sub-optimal acquisition protocols than numerical optimization methods [21]. There are, however, studies that have reported contrary results that deserve careful consideration. For example, one study has reported that neural network and random forest models trained on simulated data may produce less accurate and more biased results compared with nonlinear least-squares estimation methods [168]. Another study has shown that deep learning methods can achieve good estimation accuracy when the optimization landscape is well-behaved but that they have poor performance when the optimization landscape is degenerate such as for multi-compartment models with even two compartments [169, 153].
It has also been reported that machine learning methods are less sensitive to measurement noise and other imperfections [170, 57]. As an example, Gong et al. trained a deep learning model to compute DTI and DKI parameters using data from healthy subjects with voluntary head motion as well as attention deficit hyperactivity disorder (ADHD) patients with non-voluntary motion [57]. They found that predictions of the deep learning model were less sensitive to motion and comparable to predictions with motion-free data.
It has been shown that models that leverage spatial correlations in a neighborhood (e.g., an image patch) around the voxel of interest can lead to more accurate estimation [161, 55, 165, 171]. Early works used CNNs to learn the spatial correlations. More recent studies haves relied on attention-based neural networks that are considered to be more effective in learning correlations [162, 22, 172]. Zheng et al. argue for the superiority of transformer architectures for tissue microstructure mapping in dMRI [162]. In order to enable effective implementation of these architectures with small datasets, they introduce additional computational modules. Specifically, they use a sparse representation stage to compute the signal embeddings in a dictionary using unfolded iterative hard thresholding.
Although most studies have evaluated the new methods only in terms of estimation accuracy measures such as root mean square of the error (RMSE), several works have investigated the downstream use of the estimated tissue microstructure parameters. For example, Gibbons et al. showed that post‐stroke outcome prediction was the same for microstructure maps estimated with standard methods and those estimated with a neural network with 10-fold fewer measurements [161]. Aliotta et al. observed that deep learning-estimated FA maps resulted in more accurate delineation of the brain tumor boundaries in glioblastoma multiforme patients than with conventional methods [148]. Kaandorp et al. showed that intravoxel incoherent motion (IVIM) parameters computed by a deep learning method better predicted the chemoradiotherapy response of pancreatic cancer patients than a standard method [52]. Another work assessed the microstructure maps estimated by several deep learning methods in terms of their effectiveness in studying the impact of migraine on the brain white matter [173]. Specifically, FA and mean diffusivity (MD) values computed by different methods were compared using Tract-Based Spatial Statistics (TBSS) [174]. It was observed that, compared with a standard estimation technique, deep learning methods improved the true positive rate but also increased the false positive rate.
For DTI estimation, several works have reported good reconstructions from the theoretical minimum of six measurements (Table II). DeepDTI [46] uses six diffusion-weighted measurements and a non-weighted (b0) measurement, T1-weighted, and T2-weighted images as input and computes the residuals with respect to high-quality measurements. The estimated high-quality measurements are then used to estimate the diffusion tensor with a standard estimation technique. Detailed analysis, including local assessment of DTI-derived parameters, tract-based analysis, and tractography show that DeepDTI achieves accurate estimation while reducing the number of measurements by a factor of at least 3-5. DeepDTI is based on a 3D FCN and requires that the six diffusion-weighted measurements be acquired along the optimal directions proposed in [175]. SuperDTI, proposed by Li et al. [55], also used an FCN for direct estimation of the diffusion tensor. It achieved accurate FA computation and tractography with six measurements and accurate MD computation with three measurements. Similar results were reported in [148, 171]. Using transformer networks, another study made a similar claim and further evaluated the method using tractography and quantitative structural connectivity assessment [51]. DiffNet [149] estimated FA and MD with 3-20 diffusion-weighted measurements and showed that the estimation accuracy was still higher than standard methods. DiffNet-estimated FA was more accurate for segmenting brain tumor volume, while also reducing the scan time.
IV-D5 Other dMRI-derived parameters
In addition to the DTI, DKI, and NODDI parameters, machine learning methods have also been applied to compute other parameters. Examples include myelin water fraction [170, 42], water residence time in white matter (related to axonal permeability) [45, 176, 60], axonal radius [45, 42], spherical mean technique (SMT) [158, 168, 165], ensemble average propagator (EAP) models [177, 158], and relaxation-diffusion models of white matter microstructure [21, 178, 169]. Parik et al. developed a novel magnetic resonance fingerprinting sequence to simultaneously estimate T1-weighted, T2-weighted, and diffusion tensor parameters. Instead of the standard dictionary matching approach, they used a deep neural network to estimate the target parameters and achieved good results in healthy brains and brains of MS patients [179].
A number of works have attempted to use deep neural networks to estimate the IVIM model [163, 53, 180, 52, 181, 182, 162, 164]. One study showed that an MLP could achieve results that were comparable with or more accurate than least-squares estimation and Bayesian techniques [53]. Another study found that a neural network could outperform the state of the art estimation methods for computing the parameters of a combined IVIM-kurtosis imaging model [180]. It also showed that the neural network method made fewer predictions that were outside the range of plausible parameter values. One study showed that a neural network, trained using an unsupervised approach, produced more consistent and more accurate estimations than standard methods and it better predicted the chemoradiotherapy response of pancreatic cancer patients [52].
IV-E Fiber orientation density estimation